# Universality of Estimator for High-Dimensional Linear Models with Block Dependency

### v1

@@ A.1. Proof of Theorem 1
×$\times\left\{L_{n}^{q_{0}q_{1}}Av({L_{\ell_{i}}^{q_{1}}})\cdot\max_{Z_{ij}\in\{X_{ij},W_{ij}\}}\left(E\left[\max_{1\leq i\leq n}\sum_{j=1}^{p}|Z_{i,j}|\right]+E\left[\left(\max_{1\leq i\leq n}\sum_{j=1}^{p}|Z_{i,j}|\right)^{q_{0}q_{1}+1}\right]\right)\delta\right.$
@@
By trivial calculation, for each $\mathbf{Z}\in\{\mathbf{X},\mathbf{W}\}$, we have
@@ A.2. Proof of Theorem 3
$\leq$ $2\varepsilon_{n}+C_{1}(\max\{1,\rho_{0}^{-3}\})r_{R}(L_{n}),$
where the last two inequalities hold by assumptions. For the second term, we have
@@
where the second last inequality follows by Theorem 1. The last inequality holds from the assumption. Therefore, the required result follows. ∎
@@ A.3. Proof of Corollary 2 and Corollary 4
Whichever corollary we prove, we verify that $r_{R}(L_{n})$ converges to zero with the setup. About the minimization in terms of $\delta\in(0,\omega_{n})$, we take $\delta=\delta_{n}\searrow 0$ as $n\to\infty$ such that $L_{n}/\delta_{n}\to\infty$ holds, which implies that $\log^{2/3}_{+}(L_{n}/\delta_{n})=o(\log n)$ and $\mathcal{M}_{R}(L_{n},\delta_{n})=o(1)$ with $\delta_{n}\searrow 0$ using the convexity of $R(\cdot)$. We can find such $\delta_{n}$, since $L_{n}$ has a logarithmic rate in $n$. We set $d=o(n^{\mu}/(\log n)^{d^{*}})$ with some $\mu,d^{*}\in\mathbb{R}$ and apply the definition of $\sigma_{n}$ in (7), then obtain
@@
$=o(n^{(\bar{\mathbf{q}}_{0}+3)\alpha/{3}+2\mu/3-1/6}(\log n)^{1-d^{*}/6+(\bar{\mathbf{q}}_{0}/3+1)\kappa}).$
Therefore, we obtain $\log^{2/3}_{+}(L_{n}/\delta_{n})L_{n}^{(\bar{\mathbf{q}}_{0}+3)/{3}}\sigma^{1/3}_{n}=o(1)$, when $\mu$ is no more than $21/4-(\bar{\mathbf{q}}_{0}+3)\alpha/$ 2 and $d^{*}$ is greater than or equal to $6+(2\bar{\mathbf{q}}_{0}+6)\kappa$. Using the result, we obtain
@@
Then, at the minimiation in $\rho$, we take $\rho=\rho_{n}=\zeta_{n}^{1/7}$, then we obtain
@@ B.1. Preliminaries and Some Auxiliary Results
For the later use, we summarize the notations and the derivatives of $g(n^{-1}F_{\beta}(\cdot))$. Set $F_{\beta}(A)$ and $\bar{H}(w,A)$ as follows:
@@
$\partial_{i,j}\partial_{i,k}g(n^{-1}F_{\beta}(A))=$ $n^{-2}\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime\prime}(n^{-1}F_{\beta}(A))$
$+$ $n^{-1}g^{\prime}(n^{-1}F_{\beta}(A))\left(\langle\partial_{i,j}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\bar{H}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\right).$
where we define $\bar{H}_{\beta}:=\beta\cdot\bar{H}$ Also, we obtain the higher-order derivative as
@@
$=m^{-3}\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime\prime\prime}(n^{-1}F_{\beta}(A))$
$+$ $m^{-2}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime\prime}(n^{-1}F_{\beta}(A))\left(\langle\partial_{i,j}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right)$
$+$ $m^{-2}\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime\prime}(n^{-1}F_{\beta}(A))\left(\langle\partial_{i,k}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,k}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right)$
$+$ $m^{-2}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime\prime}(n^{-1}F_{\beta}(A))\left(\langle\partial_{i,j}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\bar{H}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\right)$
$+$ $n^{-1}g^{\prime}(n^{-1}F_{\beta}(A))$
$\times$ $\left\{\langle\partial_{i,j}\partial_{i,k}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\partial_{i,k}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,j}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right.$
$-\beta\left(\langle\partial_{i,j}\partial_{i,\ell}\bar{H}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}+\langle\partial_{i,j}\bar{H}\partial_{i,k}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\bar{H}\partial_{i,k}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right.$
$\left.+\beta\langle\partial_{i,j}\bar{H}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right)$
@@
$+\left.\beta\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\left(\langle\partial_{i,j}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right)\right\}$
@@ B.2. Proof of Lemma 5
$W_{i}^{(B_{k})}$ is defined in the similar manner of $X_{i}^{(B_{k})}$. Let $(X_{i})^{(B_{k})}_{\ell}$ denote the $\ell$-th non-zero element of $X_{i}^{(B_{k})}$. We define
@@
$+\frac{1}{2}\int^{1}_{0}(1-t)^{2}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}X_{i,\ell}X_{i,\ell^{\prime}}X_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}f(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})dt$
@@
$f(\mathbf{Z}_{i,j})$ $=f(\mathbf{Z}^{0}_{i,j})+(W^{(B_{j})}_{i})^{\top}\nabla_{i}f(\mathbf{Z}_{i,j}^{0})+\frac{1}{2}(W^{(B_{j})}_{i})^{\top}\nabla^{2}_{i}f(\mathbf{Z}_{i,j}^{0})W^{(B_{j})}_{i}$
$+\frac{1}{2}\int^{1}_{0}(1-t)^{2}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}W_{i,\ell}W_{i,\ell^{\prime}}W_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}f(\mathbf{Z}_{i,j}^{0}+tW^{(B_{j})}_{i})dt$
@@
$\nabla_{i}f(\mathbf{Z}^{0}_{i,j}):=\begin{bmatrix}\partial_{i,1}f(\mathbf{Z}^{0}_{i,j})\\ \vdots\\ \partial_{i,d}f(\mathbf{Z}^{0}_{i,j})\\ \end{bmatrix},\quad\nabla^{2}_{i}f(\mathbf{Z}^{0}_{i,j}):=\begin{bmatrix}\partial^{2}_{i,1}f(\mathbf{Z}^{0}_{i,j})&\cdots&\partial_{i,1}\partial_{i,d}f(\mathbf{Z}^{0}_{i,j})\\ \vdots&\ddots&\vdots\\ \partial_{i,1}\partial_{i,d}f(\mathbf{Z}^{0}_{i,j})&\cdots&\partial^{2}_{i,d}f(\mathbf{Z}^{0}_{i,j})\\ \end{bmatrix}.$
@@
$=$ $\sum_{i=1}^{n}\sum_{j=1}^{k}\sum_{\ell=1}^{|B_{j}|}E[((X^{(B_{j})}_{i})_{\ell}-(W^{(B_{j})}_{i})_{\ell})\partial_{i,(B_{j})_{\ell}}f(\mathbf{Z}^{0}_{i,j})]$
$+$ $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{k}\sum_{\ell=1}^{|B_{j}|}\sum_{\ell^{\prime}=1}^{|B_{j}|}E[((X^{(B_{j})}_{i})_{\ell}(X^{(B_{j})}_{i})_{\ell^{\prime}}-(W^{(B_{j})}_{i})_{\ell}(W^{(B_{j})}_{i})_{\ell^{\prime}})\partial_{i,(B_{j})_{\ell}}\partial_{,i(B_{j})_{\ell^{\prime}}}f(\mathbf{Z}^{0}_{i,j})]$
$+$ $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{k}\int^{1}_{0}(1-t)^{2}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}X_{i,\ell}X_{i,\ell^{\prime}}X_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}f(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})dt$
-- $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{k}\int^{1}_{0}(1-t)^{2}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}W_{i,\ell}W_{i,\ell^{\prime}}W_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}f(\mathbf{Z}_{i,j}^{0}+tW^{(B_{j})}_{i})dt$
By the assumption on block dependence, for any $j$ and $\ell\in[|B_{j}|]$, we have
$E[((X^{(B_{j})}_{i})_{\ell}-(W^{(B_{j})}_{i})_{\ell})\partial_{i,(B_{j})_{\ell}}f(\mathbf{Z}^{0}_{i,j})]$ $=E[E[((X^{(B_{j})}_{i})_{\ell}-(W^{(B_{j})}_{i})_{\ell})|\mathbf{Z}^{0}_{i,j}]\partial_{i,(B_{j})_{\ell}}f(\mathbf{Z}^{0}_{i,j})]$
$=E[E[((X^{(B_{j})}_{i})_{\ell}-(W^{(B_{j})}_{i})_{\ell})]\partial_{i,(B_{j})_{\ell}}f(\mathbf{Z}^{0}_{i,j})]$
@@
In the same way, or any $j$ and $\ell\in[B_{j}]$, we have
$E[((X^{(B_{j})}_{i})_{\ell}(X^{(B_{j})}_{i})_{\ell^{\prime}}-(W^{(B_{j})}_{i})_{\ell}(W^{(B_{j})}_{i})_{\ell^{\prime}})\partial_{i,(B_{j})_{\ell}}\partial_{i,(B_{j})_{\ell^{\prime}}}f(\mathbf{Z}^{0}_{i,j})]=0.$

### v2

@@ 7. Conclusion
Acknowledgments
The authors would like to thank the anonymous referees, an Associate Editor and the Editor for their constructive comments that improved the quality of this paper.
Funding
Toshiki Tsuda was supported by Grant-in-Aid for JSPS Fellows (23KJ0713). Masaaki Imaizumi was supported by JSPS KAKENHI (24K02904), JST CREST (JPMJCR21D2), JST FOREST (JPMJFR216I), and JST BOOST (JPMJBY24A9).
@@ A.1. Proof of Theorem 1
$\times\left\{L_{n}^{q_{0}q_{1}}Av({L_{\ell_{i}}^{q_{1}}})\cdot\max_{Z_{ij}\in\{X_{ij},W_{ij}\}}\left(E\left[\max_{1\leq i\leq n}\sum_{j=1}^{p}|Z_{i,j}|\right]+E\left[\left(\max_{1\leq i\leq n}\sum_{j=1}^{p}|Z_{i,j}|\right)^{q_{0}q_{1}+1}\right]\right)\delta\right$.
@@
For each $\mathbf{Z}\in\{\mathbf{X},\mathbf{W}\}$, we have
$E\left[\left(\max_{1\leq i\leq n}\sum_{j=1}^{p}|Z_{i,j}|\right)^{q_{0}q_{1}+1}\right]=$ $\left(\left\|\max_{1\leq i\leq n}\sum_{j=1}^{p}|Z_{i,j}|\right\|_{q_{0}q_{1}+1}\right)^{q_{0}q_{1}+1}$
$\leq$ $n\left(\max_{1\leq i\leq n}\left\|\sum_{j=1}^{p}|Z_{i,j}|\right\|_{q_{0}q_{1}+1}\right)^{q_{0}q_{1}+1}$
$\leq$ $n\left(\left\|\sum_{j=1}^{p}|Z_{i,j}|\right\|_{q_{0}q_{1}+1}\right)^{q_{0}q_{1}+1}$
$\leq$ $n\left(p\max_{1\leq j\leq p}\left\|Z_{i,j}\right\|_{q_{0}q_{1}+1}\right)^{q_{0}q_{1}+1},$
where the first inequality follows from Lemma 2.2.2 in van der Vaart and Wellner (2023). For the term $\max_{1\leq j\leq p}\|Z_{i,j}\|_{q_{0}q_{1}+1}$, from Assumption 3, we have
$\max_{1\leq j\leq p}\left\|Z_{i,j}\right\|_{q_{0}q_{1}+1}\lesssim_{q,M}\frac{1}{\sqrt{n}}$
Thus, we have
$E\left[\left(\max_{1\leq i\leq n}\sum_{j=1}^{p}|Z_{i,j}|\right)^{q_{0}q_{1}+1}\right]\lesssim_{\tau,q,M}n^{\frac{q_{1}q_{0}+3}{2}}.$
Hence, by $q_{0}q_{1}>0$, we also obtain
@@ A.2. Proof of Theorem 3
$\leq 2\varepsilon_{n}+C_{1}(\max\{1,\rho_{0}^{-3}\})r_{R}(L_{n}),$
where the last two inequalities hold by assumption. For the second term, we have
@@
where the second last inequality follows from Theorem 1. The last inequality holds from the assumption. Therefore, the required result follows. ∎
@@ A.3. Proof of Corollary 2 and Corollary 4
Whichever corollary we prove, we verify that $r_{R}(L_{n})$ converges to zero with the setup. For the minimization in terms of $\delta\in(0,\omega_{n})$, we take $\delta=\delta_{n}\asymp n^{-c_{1}}$ as $n\to\infty$ with some $c_{1}>1$. Since $L_{n}$ has an order $O(n\log^{\kappa}n)$, we have $\log^{2/3}_{+}(L_{n}/\delta_{n})=O(\log^{2/3}n)=o(\log n)$. In addition, we have $\mathcal{M}_{R}(L_{n},\delta_{n})=o(1)$ with $\delta_{n}\searrow 0$. Here, we apply the locally Lipschitz or pseudo-Lipschitz property of the specified regularizer to state this result. We set $d=o(n^{\mu}/(\log n)^{d^{*}})$ with some $\mu,d^{*}\in\mathbb{R}$ and apply the definition of $\sigma_{n}$ in (8), then obtain
@@
$=o(n^{(\bar{\mathbf{q}}_{0}+3)\alpha/{3}+2\mu/3-1/6}(\log n)^{1-2d^{*}/3+(\bar{\mathbf{q}}_{0}/3+1)\kappa}).$
Therefore, we obtain $\log^{2/3}_{+}(L_{n}/\delta_{n})L_{n}^{(\bar{\mathbf{q}}_{0}+3)/{3}}\sigma^{1/3}_{n}=o(1)$, when $\mu$ is no more than $1/4-(\bar{\mathbf{q}}_{0}+3)\alpha/2$ and $d^{*}$ is greater than or equal to $3/2+1/2(\bar{\mathbf{q}}_{0}+3)\kappa$. Using the result, we obtain
@@
Then, at the minimization in $\rho$, we take $\rho=\rho_{n}=\zeta_{n}^{1/7}$, then we obtain
@@ Appendix B Proof of lemmas and propositions
In this section, we provide several proofs. We remind some notation. For a matrix $A\in\mathbb{R}^{d_{1}\times d_{2}}$ and a function $h:\mathbb{R}^{d_{1}\times d_{2}}\to\mathbb{R}$, $\partial_{j_{1},j_{2}}h(A):=\partial h(A)/\partial{A_{j_{1},j_{2}}}$ denotes a partial derivative in terms of a $(j_{1},j_{2})$-element $A_{j_{1},j_{2}}$ of $A$.
@@ B.1. Preliminaries and some auxiliary results
For later use, we summarize the notations and the derivatives of $g(n^{-1}F_{\beta}(\cdot))$. Set $F_{\beta}(A)$ and $\bar{H}(w,A)$ as follows:
@@
$\partial_{i,j}\partial_{i,k}g(n^{-1}F_{\beta}(A))$
$=n^{-2}\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime\prime}(n^{-1}F_{\beta}(A))$
$\quad+n^{-1}g^{\prime}(n^{-1}F_{\beta}(A))\left(\langle\partial_{i,j}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\bar{H}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\right).$
where we define $\bar{H}_{\beta}:=\beta\cdot\bar{H}$ . Also, we obtain the higher-order derivative as
@@
$=n^{-3}\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime\prime\prime}(n^{-1}F_{\beta}(A))$
$~+n^{-2}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime\prime}(n^{-1}F_{\beta}(A))\left(\langle\partial_{i,j}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right)$
$~+n^{-2}\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime\prime}(n^{-1}F_{\beta}(A))\left(\langle\partial_{i,k}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,k}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right)$
$~+n^{-2}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime\prime}(n^{-1}F_{\beta}(A))\left(\langle\partial_{i,j}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\bar{H}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\right)$
$~+n^{-1}g^{\prime}(n^{-1}F_{\beta}(A))$
$~\times\left\{\langle\partial_{i,j}\partial_{i,k}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\partial_{i,k}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,j}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right$.
$~-\beta\left(\langle\partial_{i,j}\partial_{i,\ell}\bar{H}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}+\langle\partial_{i,j}\bar{H}\partial_{i,k}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\bar{H}\partial_{i,k}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right$.
$~\left.+\beta\langle\partial_{i,j}\bar{H}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right)$
@@
$~+\left.\beta\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\left(\langle\partial_{i,j}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right)\right\}.$
@@ B.2. Proof of Lemma 5
$W_{i}^{(B_{k})}$ is defined in the similar manner of $X_{i}^{(B_{k})}$. We define
@@
$~+\frac{1}{2}\int^{1}_{0}(1-t)^{2}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}X_{i,\ell}X_{i,\ell^{\prime}}X_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}f(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})dt,$
@@
$f(\mathbf{Z}_{i,j-1})$ $=f(\mathbf{Z}^{0}_{i,j})+(W^{(B_{j})}_{i})^{\top}\nabla_{i}f(\mathbf{Z}_{i,j}^{0})+\frac{1}{2}(W^{(B_{j})}_{i})^{\top}\nabla^{2}_{i}f(\mathbf{Z}_{i,j}^{0})W^{(B_{j})}_{i}$
$~+\frac{1}{2}\int^{1}_{0}(1-t)^{2}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}W_{i,\ell}W_{i,\ell^{\prime}}W_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}f(\mathbf{Z}_{i,j}^{0}+tW^{(B_{j})}_{i})dt,$
@@
$\nabla_{i}f(\mathbf{Z}^{0}_{i,j}):=\begin{bmatrix}\partial_{i,1}f(\mathbf{Z}^{0}_{i,j})\\ \vdots\\ \partial_{i,p}f(\mathbf{Z}^{0}_{i,j})\\ \end{bmatrix},\quad\nabla^{2}_{i}f(\mathbf{Z}^{0}_{i,j}):=\begin{bmatrix}\partial^{2}_{i,1}f(\mathbf{Z}^{0}_{i,j})&\cdots&\partial_{i,1}\partial_{i,p}f(\mathbf{Z}^{0}_{i,j})\\ \vdots&\ddots&\vdots\\ \partial_{i,1}\partial_{i,p}f(\mathbf{Z}^{0}_{i,j})&\cdots&\partial^{2}_{i,p}f(\mathbf{Z}^{0}_{i,j})\\ \end{bmatrix}$.
@@
$=\sum_{i=1}^{n}\sum_{j=1}^{k}\sum_{\ell\in B_{j}}E[(X_{i,\ell}-W_{i,\ell})\partial_{i,\ell}f(\mathbf{Z}^{0}_{i,j})]$
$~+\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{k}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}E[(X_{i,\ell}X_{i,\ell^{\prime}}-W_{i,\ell}W_{i,\ell^{\prime}})\partial_{i,\ell}\partial_{i,\ell^{\prime}}f(\mathbf{Z}^{0}_{i,j})]$
$~+\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{k}\int^{1}_{0}(1-t)^{2}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}X_{i,\ell}X_{i,\ell^{\prime}}X_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}f(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})dt$
$~-\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{k}\int^{1}_{0}(1-t)^{2}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}W_{i,\ell}W_{i,\ell^{\prime}}W_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}f(\mathbf{Z}_{i,j}^{0}+tW^{(B_{j})}_{i})dt.$
By the assumption on block dependence, for any $j$ and $\ell\in B_{j}$, we have
$E[(X_{i,\ell}-W_{i,\ell})\partial_{i,\ell}f(\mathbf{Z}^{0}_{i,j})]$ $=E[E[(X_{i,\ell}-W_{i,\ell})|\mathbf{Z}^{0}_{i,j}]\partial_{i,\ell}f(\mathbf{Z}^{0}_{i,j})]$
$=E[E[(X_{i,\ell}-W_{i,\ell})]\partial_{i,\ell}f(\mathbf{Z}^{0}_{i,j})]$
@@
In the same way, for any $j$ and $\ell,\ell^{\prime}\in B_{j}$, we have
$E[(X_{i,\ell}X_{i,\ell^{\prime}}-W_{i,\ell}W_{i,\ell^{\prime}})\partial_{i,\ell}\partial_{i,\ell^{\prime}}f(\mathbf{Z}^{0}_{i,j})]=0.$